OpenStudy (ashy98):

Complete the table below to solve the equation 2.5x – 10.5 = 64(0.5x). x f(x) = 2.5x – 10.5 g(x) = 64(0.5x) 2 3 4 5 6

OpenStudy (ganpat):

substitute value of x as 2,3,4,5 & 6 in given equation.. x f(x) = 2.5x – 10.5 g(x) = 64(0.5x) 2 2.5 (2)- 10.5 64(0.5 * 2) 3 2.5(3) - 10.5 64(0.5 *3)

OpenStudy (haseeb96):

just put x=2 in the f(x) and g(x) functions u can easily fill the whole table

OpenStudy (haseeb96):

@Ashy98 Did u get my point?

OpenStudy (haseeb96):

i m going to give u an example here that is if f(x) = x^3+3 and g(x)= x^4+2 now solve the table by 2 and 3 what would i do i will take first x=2 then i will put it in the both equation f(2)= (2)^3 +3 and g(2)=(2)^4 +2 = 8+3 and =16+2 = 11 and =18 by this way u can easily solve the equation

OpenStudy (ashy98):

x f(x) = 2.5x 0.5 g(x) = 64(0.5x) 2 2.5(2)- 10.5 64(0.5 * 2) 3 2.5(3) - 10.5 64(0.5 *3) 4 2.5(4) - 10.5 64(0.5 *4) 5 2.5(5) - 10.5 64(0.5 *5) 6 2.5(6) - 10.5 64(0.5 *6) like that?

OpenStudy (haseeb96):

yes in this way but u should solve it more like i did

OpenStudy (ganpat):

yeah @Ashy98 .. I just explained you the method.. You have to simplify as @Haseeb96 did..

OpenStudy (haseeb96):

Because in question it is asking that u have to solve the equation \and we know what its means.

OpenStudy (ashy98):

okay thanks guys!

OpenStudy (haseeb96):

@Ashy98 Most Welcome lady